AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 16 of 33  ELOG logo
ID Date Author Subject Text Attachments
  359   Tue Mar 1 23:43:15 2022 David, GeorgeBeam Intensity Drop00:38
Intensity in the ring dropped by 2
orders of magnitude to 5e5 on injection
  
  358   Tue Mar 1 23:30:04 2022 David, GeorgeCurrent status00:30
Pressure in S1 slightly decreased
from before but still high enough to measure
 2022-03-02_00-39-47xytotal.png2022-03-02_00-39-20exRate.png2022-03-02_00-38-44extotal.png 
  357   Tue Mar 1 22:10:26 2022 David, GeorgeCurrent statusRing current 6.67e7 particles/s

pressure at S1: 2.3-2.8e-8

Merger, Tape server (WS3):
 2022-03-01_23-14-43aida.png2022-03-01_23-13-57exrate.png2022-03-01_23-13-40xyto.png2022-03-01_23-13-10extot.png 
  356   Tue Mar 1 21:34:06 2022 PM, LVcheckno change:

Check Clock Status': Passed
5, Failed 0
 14x 
  355   Tue Mar 1 20:41:28 2022 PM, LVcheckno change:

Check Clock Status': Passed
5, Failed 0
 14x 
  354   Tue Mar 1 19:41:24 2022 PM, LVtarget temperature walkthroughA target temperature walkthrough has been
initiated at 20:41 since the target density
at S1 had slightly decreased (with ~0.1-0.2)
  
  353   Tue Mar 1 19:40:28 2022 PM, LVcheckno change:

Check Clock Status': Passed
5, Failed 0
 12x 
  352   Tue Mar 1 18:43:41 2022 PM, LVcheckno change:

Check Clock Status': Passed
5, Failed 0
 14x 
  351   Tue Mar 1 17:36:04 2022 PM, LVcheckno change:

Check Clock Status': Passed
5, Failed 0
 14x 
  350   Tue Mar 1 16:57:34 2022 CB, LV, PMHNew runBeam issues due to too long storage time
in the last hours. Motors moved out and not
moving.
  
  349   Tue Mar 1 14:01:01 2022 MS, SHBcheck'Check Clock Status', 'Check
ADC Calibration' and 'Check the White
rabbit decoder status' : no change
 11x 
  348   Tue Mar 1 12:58:32 2022 CB, MS, SS, MLSynced pneumatic motionNew run R84

Attempting to sync pneumatic movement
with ring. Still see something on t out ->
  
  347   Tue Mar 1 12:33:21 2022 MS, SHBcheck'Check Clock Status', 'Check
ADC Calibration' and 'Check the White
rabbit decoder status': no change
 11x 
  346   Tue Mar 1 10:43:36 2022 MS, SHBcheck 'Check Clock Status', 'Check
ADC Calibration' and 'Check the White
rabbit decoder status': same
 11x 
  345   Tue Mar 1 10:35:59 2022 MS, SHBnext run on tape server tabold potentiometer position 19.1 mm

new run number: 83

new potentiometer position : 27.58
  
  344   Tue Mar 1 08:02:26 2022 SHB, MScheck 

'Check Clock Status', 'Check
ADC Calibration' and 'Check the White
 11x 
  343   Tue Mar 1 06:35:42 2022 SHB, MScheckdetector motors are still out

'Check Clock Status', 'Check
ADC Calibration' : Clock status test
 11x 
  342   Tue Mar 1 06:04:04 2022 SHB, MSpositionsjust FYI: current positions saved in files
are x=1000 and y=-750. this information is
good in case of crash.
  
  341   Tue Mar 1 05:23:02 2022 JM, EMCurrent status 06:20

Servo motors moved to fully out
position as some work on the beam was neccessary.
 2022-03-01_06-42-53-ExRate.png 
  340   Tue Mar 1 04:07:31 2022 JM, EMTemperature cycleHave been unable to achieve decent target
density with temperature cycling so far.
Will attempt to go up to 300 k and back down
  
ELOG V3.1.4-unknown